A major component of gasoline is octane (C8H18). When octane is burned in air, it chemically reacts with oxygen gas (O2) to produce carbon dioxide (CO2) and water (H2O).
What mass of octane is consumed by the reaction of 2.3 g of oxygen gas?
1
Expert's answer
2020-09-14T08:26:33-0400

At first chemical reaction should be written and balanced:

2C8H18 + 25O2 = 16CO2 + 18H2O.

The molar masses of reactants should be found:

M(O2)=2×16=32(g/mol),

M(C8H18)=8×12+18×1=114(g/mol).

Taking into account the reaction coefficients the mass of octane is found from a proportion:

x=2.3×2×114/25×32=0.6555(g).

Answer: 0.6555 g of octane.